OpenStudy (anonymous):

6,561^(2n)=729^(n+2)

OpenStudy (anonymous):

Take logs of both sides 2nlog6561 = (n+2)log729, then solve for n

myininaya (myininaya):

3^(8*2n)=3^(6(n+2))

myininaya (myininaya):

16n=6n+12 16n-6n=12 10n=12 n=12/10=6/5

myininaya (myininaya):

hey and gj on those parathesis dvinyard :)

OpenStudy (anonymous):

thanks haha